Diogo has a utility function , where A, α and β are constants, B is burritos, and Z is pizzas. If the price of burritos, Pb, is N$2 and the price of pizzas, Pz is N$1, and Y is N$100, what is Diogo’s optimal bundle?
1
Expert's answer
2015-05-02T06:52:36-0400
If Diogo has a utility function U(B, Z) = AB^αZ^β and if the price of burritos, Pb, is N$2 and the price of pizzas, Pz is N$1, and Y is N$100, using the Lagrangian method we can find the optimal bundle:
Qb = β/(α + β)*100/2 = 50β/(α + β)
Qz = 100β/(α + β)